# Are LLMs Smarter Than Chimpanzees? An Evaluation on Perspective Taking and Knowledge State Estimation

### v1

@@ 5.2 Further Analysis and Ablation
We examine whether narrative style affects performance. We compare five genres in our benchmark dataset: romance, fantasy, science fiction, kids, and mystery. As shown in Figure 3, different genres show no clear performance difference across either task. For logic error detection, the fantasy genre shows slightly lower performance—possibly because models tend to excuse errors as explainable within the fantastic setting. [^5] The stories explicitly state that the characters in our questions do not possess magic power (so they cannot use magic to obtain implausible knowledge).
@@
The results are shown in Figure 4. On average, the explicit prompt to pay attention to character knowledge improves accuracy by 12.6 pp in IKD (Whole Set) and 12.2 pp in KNP, with DeepSeek-R1 achieving the highest boost. This suggests the LLMs do not yet understand the dependency between actions taken by humans and their knowledge, which is a critical component of USI according to Tomasello et al. (2005). After being prompted, the best accuracy is 84% on KNP, still falling short of the performance of the human participants who did not receive this prompt.
@@
Qwen3-8B (think) 0.08 0.12 0.14 0.33
Qwen3-32B (think) 0.07 0.34 0.29 0.60
Qwen3-Max 0.13 0.39 0.36 0.72
@@
LLaMA3.1-70B 0.08 0.34 0.38 0.63
LLaMA3.3-70B 0.05 0.26 0.30 0.58
Kimi-K2-Instruct 0.19 0.52 0.59 0.76
Kimi-K2-Thinking 0.37 0.60 0.61 0.74
DeepSeek-V3.2 0.22 0.35 0.41 0.79
DeepSeek-R1-V3.2 0.19 0.26 0.29 0.64
GPT-4o 0.03 0.16 0.21 0.61
GPT-5 (high) 0.14 0.20 0.19 0.47
Gemini-3-Pro 0.48 0.67 0.71 0.80
Claude-4.5-Opus 0.31 0.46 0.45 0.61
@@
Qwen3-8B (think) 0.56 0.59
Qwen3-32B (think) 0.55 0.60
Qwen3-Max 0.60 0.57
@@
LLaMA3.1-70B 0.54 0.56
LLaMA3.3-70B 0.53 0.59
Kimi-K2-Instruct 0.55 0.53
Kimi-K2-Thinking 0.56 0.54
DeepSeek-V3.2 0.55 0.53
DeepSeek-R1-V3.2 0.57 0.59
GPT-4o 0.51 0.51
GPT-5 (high) 0.63 0.59
Gemini-3-Pro 0.70 0.67
Claude-4.5-Opus 0.71 0.66
@@ 5.3 Case Studies
Prediction results on the IKD task, shown in Figures 5-7, indicate that current LLMs are not very sensitive to implausible knowledge. Although Claude-4.5-Opus and Gemini-3-Pro perform better than others, they still show a gap compared to humans. Gemini-3-Pro, specifically, shows a tendency to be overly strict and tries to locate errors even in correct stories (Figure 7).
@@ 6 Conclusion
In this work, we explore whether current LLMs possess a significant Theory-of-Mind (ToM) capacity that distinguishes human intelligence from other animals: understanding others’ independent knowledge states and predicting their corresponding actions. We carefully design two tasks for evaluation: Implausible Knowledge Detection (IKD) and Knowledge-sensitive Next-Action Prediction (KNP). Based on experimental exploration using our proposed benchmarks, most leading LLMs show near-random performance in this area, with a significant gap compared to humans. This finding highlights the need for future research.
@@ A.2 Automatic Data Construction Prompts
The prompt for an erroneous story construction (IKD task) is provided in Table 5. The prompt for Next-action Question Construction (KNP task) is provided in Table 6.
@@ A.3 Human Refinement
KNP Task:
Annotators refined approximately 40% of the auto-generated data. Our authors then reviewed all refined data, making additional improvements to 8% of the data. We will release the data annotation file (containing detailed instructions and annotation samples) to support future research.
@@ Appendix B Evaluation Prompts
For tested LLMs, we apply the prompt shown in Table 8 for the IKD task and the one shown in Table 9 for the KNP task.
@@
Erroneous Story Construction
Original story:
<original_story>
{story_original}
</original_story>
Instructions:
1. Introduce implausible knowledge for a character:
 - The knowledge should be mentioned in the original story.
Notes:
- Ensure the rewritten event is fluent and coherent.
Output Format:
<response>
<original_event>
</original_event>
<character>
[the name of the character]
</character>
<implausible_knowledge>
</implausible_knowledge>
<knowledge_explanation>
[explain and verify that the character cannot know this knowledge]
</knowledge_explanation>
<modified_explanation>
</modified_explanation>
<modified_event>
</modified_event>
<modified_story>
</modified_story>
</response>
Table 5: The prompt for generating erroneous stories.
@@ 3. Construct the answers: Provide two answers:
KNP task
Read the story below and answer the related question.
Story:
{story}
Question:
{question}
Answers:
{answers}
Output Format:
Reasoning:
{The reasoning process}
Answer: {“A” or “B”}
Table 9: The prompt for KNP Task.

### v2

@@ 5.2 Further Analysis and Ablation
We examine whether narrative style affects performance. We compare five genres in our benchmark dataset: romance, fantasy, science fiction, kids, and mystery. As shown in Figure 3, different genres show no clear performance difference across either task. The fantasy and mystery genres show slightly lower performance—possibly because models tend to excuse errors as explainable within the fantastic and some mysterious settings. [^6] The stories explicitly state that the characters in our questions do not possess magic power (so they cannot use magic to obtain implausible knowledge).
@@
The results are shown in Figure 4. On average, the explicit prompt to pay attention to character knowledge improves accuracy by 12.6 pp in IKD (Whole Set) and 14.8 pp in KNP, with GPT-5.4 achieving the highest boost. This suggests the LLMs do not yet understand the dependency between actions taken by humans and their knowledge, which is a critical component of USI according to Tomasello et al. (2005). After being prompted, the best accuracy is 85% on KNP, still falling short of the performance of the human participants who did not receive this prompt.
@@
Qwen3-8B (think) 0.05 0.12 0.14 0.33
Qwen3-32B (think) 0.10 0.34 0.29 0.60
Qwen3-Max (think) 0.07 0.28 0.33 0.69
@@
LLaMA3.1-70B 0.08 0.29 0.39 0.67
LLaMA3.3-70B 0.05 0.18 0.3 0.55
Kimi-K2.5-Instruct 0.27 0.47 0.54 0.70
Kimi-K2.5-Thinking 0.26 0.46 0.54 0.71
DeepSeek-V3.2 0.20 0.36 0.41 0.78
DeepSeek-R1-V3.2 0.16 0.26 0.30 0.59
GPT-4o 0.03 0.12 0.22 0.62
GPT-5.4 (high) 0.12 0.22 0.30 0.61
Gemini-3-Pro 0.47 0.68 0.73 0.80
Claude-4.5-Opus 0.30 0.45 0.48 0.62
@@
Qwen3-8B (think) 0.49 0.53
Qwen3-32B (think) 0.49 0.51
Qwen3-Max (think) 0.53 0.53
@@
LLaMA3.1-70B 0.55 0.57
LLaMA3.3-70B 0.54 0.58
Kimi-K2.5-Instruct 0.56 0.56
Kimi-K2.5-Thinking 0.57 0.56
DeepSeek-V3.2 0.49 0.51
DeepSeek-R1-V3.2 0.57 0.58
GPT-4o 0.49 0.52
GPT-5.4 (high) 0.61 0.62
Gemini-3-Pro 0.67 0.64
Claude-4.5-Opus 0.69 0.64
@@ 5.3 Case Studies
Prediction results on the IKD task, shown in Figures 5-7, indicate that current LLMs are not very sensitive to actions driven by implausible knowledge when processing stories. Although Claude-4.5-Opus and Gemini-3-Pro perform better than others, they still show a gap compared to humans. Gemini-3-Pro, specifically, shows a tendency to be overly strict and tries to locate errors even in correct stories (Figure 7).
@@ 6 Conclusion
In this work, we explore whether current LLMs possess a significant Theory-of-Mind (ToM) capacity that distinguishes human intelligence from other animals: understanding others’ independent knowledge states and predicting their corresponding actions. We carefully design two tasks for evaluation: Knowledge-sensitive Next-Action Prediction (KNP) and Implausible Knowledge Detection (IKD). Based on experimental exploration using our proposed benchmarks, most leading LLMs show near-random performance in this area, with a significant gap compared to humans. This finding highlights the need for future research.
@@ A.2 Automatic Data Construction Prompts
The prompt for Next-action Question Construction (KNP task) is provided in Table 5. The prompt for an erroneous story construction (IKD task) is provided in Table 6.
@@ A.3 Human Refinement
Annotators refined approximately 40% of the auto-generated data. Our authors then reviewed all refined data, making additional improvements to the data. We will release the data annotation file (containing detailed instructions and annotation samples) to support future research.
@@ Appendix B Evaluation Prompts
For tested LLMs, we apply the prompt shown in Table 8 for the KNP task and the one shown in Table 9 for the IKD task.
@@ 3. Construct the answers: Provide two answers:
Erroneous Story Construction
Original story:
<original_story>
{story_original}
</original_story>
Instructions:
1. Introduce implausible knowledge for a character:
 - The knowledge should be mentioned in the original story.
Notes:
- Ensure the rewritten event is fluent and coherent.
Output Format:
<response>
<original_event>
</original_event>
<character>
[the name of the character]
</character>
<implausible_knowledge>
</implausible_knowledge>
<knowledge_explanation>
[explain and verify that the character cannot know this knowledge]
</knowledge_explanation>
<modified_explanation>
</modified_explanation>
<modified_event>
</modified_event>
<modified_story>
</modified_story>
</response>
Table 6: The prompt for generating erroneous stories.
@@ 1. Introduce implausible knowledge for a character:
KNP task
Read the story below and answer the related question.
Story:
{story}
Question:
{question}
Answers:
{answers}
Output Format:
Reasoning:
{The reasoning process}
Answer: {“A” or “B”}
Table 8: The prompt for KNP Task.